OpenStudy (anonymous):

The perimeter of an isosceles right angled triangle is 2p cm. Find out the area of the same triangle. (A) (2 + root 2 )p^2 cm2 (B) (2 - root– 2 )p^2 cm2 (C) (3- 2 root 2 )p^2 cm2 (D) (3 + 2 root 2 )p^2 cm2

OpenStudy (amistre64):

iso rt tri; isnt that a 45-45-90?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

1,1,sqrt(2)

OpenStudy (amistre64):

x(1+1+sqrt(2)) = 2pi maybe?

OpenStudy (amistre64):

mistook p tor pi lol

OpenStudy (anonymous):

so s + s + sroot2 = 2p find s A = 0.5 s^2

OpenStudy (anonymous):

Are we to substitute s=2p/(2+ root2)?

OpenStudy (amistre64):

was trying to find the scalar used to make the perimeter equal 2pi to adjust the legs :) good idea, wrong problem tho lol

OpenStudy (anonymous):

So, I am getting (2p^2)/6+4 root 2 as the final answer(after simplication). But exactly which option is that expression matching? If any mistake is there, please find out.

OpenStudy (amistre64):

s(2 + sqrt(2)) = 2p s = 2p/(2+sqrt(2)) [2p/(2+sqrt(2))]^2 4p^2 -------------- = -------------- 2/1 2(4+2+4sqrt(2)) 2p^2 p^2 -------------- = ----------- ; if I did it right 6+4sqrt(2) 3+2sqrt(2)

OpenStudy (anonymous):

So, which option is correct?

OpenStudy (amistre64):

dunno yet, gotta see where i made a mistake first lol

OpenStudy (anonymous):

I think u have solved correctly. I think the step may be to rationalise the expression.

OpenStudy (amistre64):

good thinking

OpenStudy (amistre64):

p^2 3-2sqrt(2) p^2 (3-2sqrt(2)) ----------- * ---------- = -------------- 3+2sqrt(2) 3-2sqrt(2) 9-4(2)

OpenStudy (amistre64):

C is my best bet

OpenStudy (anonymous):

Eureka!!!! you're right. It can be rewritten as (3-2 root 2)(p^2).So, obviously option(c) is correct. Thanks a lot..

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

Thanks. If any doubt, I will come tomorrow mostly at the same time with several maths questions. I hope u answer them. Thanks a lot!!! Gudbye!!!